Circular No. 6612

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bn IN UGC 4329 Qiyuan Qiao, Weidong Li, Yulei Qiu, Yi Zhang, and Jingyao Hu, Bei-

jing Astronomical Observatory (BAO), on behalf of the BAO Supernova Survey, report their discovery of a supernova in UGC 4329, found on unfiltered CCD images taken on Apr. 3.6 and 4.5 UT with the BAO 0.6-m reflector. The supernova brightened from mag 16.4 to 16.2 between the two dates. SN 1997bn is located at α = 8^{h}19^{m}02^{s}.3, δ = +21^{o}11′01′′ (equinox 2000.0), which is 6′′.0 east and 8′′.2 south of the galaxy’s nucleus. Unfiltered CCD images of the same field taken on Mar. 27.6 (limiting mag ∼ 18.0) show no star at the position of the supernova. A low-resolution spectrum taken by Y.-q. Qu, Chinese University of Science and Technology; and J.- y. Wei, D.-w. Xu, and W.-d. Li, BAO, on Apr. 4.7 with the BAO 2.16-m telescope of Xinglong station shows this to be a type-II supernova prior to optical maximum. Weak yet conspicuous hydrogen Balmer lines are seen superimposed on a very blue continuum. The He I 587.6-nm line is also present. The expansion velocity derived from the P-Cyg profile of Hα is ∼ 16 000 km/s. The recession velocity of the host galaxy is ∼ 6300 km/s, as derived from the narrow Hα emission line.

SUPERNOVA 1997bo IN ANONYMOUS GALAXY The BAO Supernova Survey also reports their discovery of another

supernova (mag ∼ 16.1) on unfiltered CCD images taken on Apr. 3.6 and 4.6 UT with the BAO 0.6-m reflector. SN 1997bo is located at α = 12^{h}19^{m}52^{s}.6, δ = +7^{o}43′50′′ (equinox 2000.0). Prediscovery images yield the following magnitudes: Mar. 15, 15.9; 27, 16.1. No star was visible on an image taken on 1991 Jan. 27. The supernova is several times brighter than the host galaxy. A nearby star (R ∼ 15.4) has position end figures 53^{s}.01, 48′′.3. A low-resolution spectrum, taken on Apr. 4.8 by J.-y. Wei, D.-w. Xu, and W.-d. Li, and Y.-q. Qu, shows this to be a type-II supernova, possibly at plateau phase. There are very strong hydrogen Balmer lines, and the Ca II infrared triplet is also present. There are some Fe lines at the blue end of the spectrum, though it is difficult to discern them because of the medium signal-to-noise ratio. The expansion velocity of the supernova envelope is ∼ 7000 km/s, as derived from the absorption minimum of Hα. There is strong, narrow Hα and O I 557.7-nm emission of the host galaxy on the spectrum, with the recession velocity measured as ∼ 3690 km/s.

1997 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green